# Mind the Gap: Aligning Knowledge Bases with User Needs to Enhance Mental Health Retrieval

### v1

@@ 1 Introduction
Access to reliable and timely mental health information is essential for early help-seeking and psychological support (Pretorius et al., 2019). Kalckreuth et al. (2014) found that 70.9% of psychiatric patients used the Internet for mental health purposes, most often to learn about disorders, medications, or services. With the growing reliance on self-help platforms, forums, chatbots, and apps (He et al., 2019), retrieving accurate and trusted psychoeducational resources has become increasingly important–both when accessed directly or when surfaced in chatbot responses through retrieval-augmented generation (RAG) (Kampman et al., 2024).
@@
RAG is a natural language processing approach that bridges fluent generation and factual grounding by combining the semantic strengths of large language models (LLMs) with evidence from trusted knowledge bases (Oche et al., 2025; Lewis et al., 2020). By conditioning outputs on retrieved documents rather than relying solely on pretrained parameters, RAG reduces hallucinations and produces more accurate, contextually appropriate, and trustworthy responses (Ayala and Bechard, 2024; Chung et al., 2023; Sharma, 2025; He et al., 2025).
@@
In mental health, knowledge bases that power retrieval systems usually contain articles, exercises, and therapeutic guides curated by trained professionals. These are often national platforms tailored to local contexts (e.g., seasonal affective disorder is absent in tropical countries like Singapore), but they are not designed for automated retrieval and are costly to expand–especially with evolving user needs. As a result, they often lack coverage across the full spectrum of user needs, both in content and in alignment with user style, language, and context (Ibrahim et al., 2021; Mao et al., 2022; He et al., 2022). For instance, a systematic review inventory by the Swedish Agency for Health Technology Assessment (SBU) identified over 2,000 evidence gaps in mental health from 2005–2020, highlighting severe content gaps where no systematic review exists or where evidence remains inconclusive (Swedish Agency for Health Technology Assessment and Assessment of Social Services (2021), SBU). We posit that one powerful way of characterizing such gaps is to overlay the knowledge base with an organic source of expression of user needs: gaps occur when topics are frequently raised by users in queries, surveys, or helpline conversations but are underrepresented or absent in the corpus. Ultimately, the effectiveness of information retrieval depends on the quality and completeness of the underlying corpus (Li et al., 2024; He et al., 2017).
@@
Recent studies have examined how gaps between user demand and document supply affect retrieval and content generation. Abián et al. (2022) showed that such misalignments can be systematically identified, but their coverage analysis relies on Wikipedia pageview metrics and lacks generalizability beyond that domain. Sinha et al. (2023) introduced the metric of findability, which measures how easily relevant documents can be surfaced by real queries even if buried in a corpus, distinguishing it from traditional retrievability metrics. However, findability remains a passive diagnostic that does not actively resolve coverage gaps. In contrast, Kang et al. (2025) proposed CCQGen, a concept-coverage–based query generation framework that iteratively produces synthetic queries to expose uncovered concepts within individual documents. While CCQGen improves internal document representation, it does not incorporate real user demand or address under-served topics at the corpus level.
@@
Our approach differs from prior work by explicitly integrating demand-driven gap detection with adaptive synthetic content generation, allowing us to identify under-served areas of the corpus based on real user demand and strategically generate targeted content to close those gaps. In this paper, we take a systematic and automated approach to improving RAG methods at the source: finding and filling content gaps rather than only optimizing retrieval or ranking. To demonstrate the utility of our framework, we apply it to a large mental health knowledge base. Our contributions are threefold:
@@
  • Knowledge base gap analysis using real-world conversations. We present a systematic framework for identifying topic-level content gaps. Using the mental health resources from mindline.sg (Weng et al., 2024), Singapore’s national digital mental health platform, as a case study, we show how real users’ posts on the let’s talk anonymous mental health forum can reveal underserved needs.
@@ 2.1 Data Collection and Preprocessing
We used two primary data sources: a mental health corpus of psychoeducational resources and a real-world source of expressed user concerns. The knowledge base consisted of 387 curated documents from mindline.sg (Weng et al., 2024). The naturalistically expressed user queries were based on 1,223 publicly available, anonymous posts in the ”Ask a Therapist” subforum of let’s talk where users can pose questions for professional counselors, thus indicating individual needs (MOH Office for Healthcare Transformation, 2025). The posts were split into 978 (80%) training queries for gap identification and augmentation and 245 (20%) test queries for evaluation.
@@ 2.1.1 Topic Taxonomy and Annotation
Resources and user queries were categorized into the same list of topics to enable matching. We based our topic taxonomy on the Clinician Index of Client Concerns (CLICC), a framework used by mental health professionals to categorize concerns presented by clients (for Collegiate Mental Health, 2019). To improve interpretability of these 46 CLICC topics, we expanded each topic into eight more granular subtopics, refined with therapist input, yielding 368 clinically relevant categories (see Appendix A). Using GPT-4o (see Appendix B for prompt), we labeled the 978 user queries and 387 mindline.sg resources with a single representative topic and subtopic.
@@ 2.2 Content Gap Analysis
Our proposal is that knowledge base gaps can be identified by overlaying naturalistic data indicating demand for certain content and topics. We demonstrate this with posts from a mental health forum. We applied a gap analysis framework to the mindline.sg corpus, assigning each query and document a subtopic for topic-level scoring. We defined two metrics, coverage gap and usefulness gap, which were computed and combined into an equal-weight hybrid score.
@@ 2.2.1 Coverage Gap: TF-IDF–Inspired Scoring
This metric is inspired by the Term Frequency-Inverse Document Frequency (TF-IDF) formula (Salton and Buckley, 1988), which measures how important a word is to a particular document in a collection by weighing a term’s local frequency against its global rarity:
@@
We adapted the TF–IDF concept from the word–document setting to the subtopic–query setting. In standard TF–IDF, a word is considered important if it occurs frequently in a given document (local frequency) but is relatively rare across the entire corpus (global rarity). Analogously, in our formulation, a subtopic is considered under-served if it appears frequently in user queries (local demand) but is covered by relatively few documents in the knowledge base (global rarity). This adaptation allows us to quantify coverage gaps as the imbalance between what users ask about and what resources actually exist.
@@
where $f_{p}(t)$ is the number of queries labeled with subtopic $t$, $df(t)$ is the number of mindline.sg documents covering $t$, $D$ is the total number of documents, $c=1$ is a smoothing constant, and $\alpha=1.5$ amplifies the effect of underrepresented topics.
Thus, the score highlights subtopics with high demand but low supply, identifying knowledge gaps that limit the users’ ability to find needed resources.
@@ 2.2.2 Usefulness Gap: LLM-as-a-Judge
We adopt the LLM-as-a-Judge paradigm, where LLMs evaluate content quality. Prior work shows that models like GPT-4 can reach about 80% agreement with human preferences, a level similar to agreement between humans, when tested on both expert-designed questions and large-scale crowdsourced comparisons (Zheng et al., 2023). This demonstrates the scalability of LLM-as-a-Judge, though limitations such as bias and reasoning gaps remain. Recent frameworks extend this approach to retrieval-augmented generation (RAG), focusing on Context Relevance, Answer Faithfulness, and Answer Relevance (Es et al., 2024; Saad-Falcon et al., 2024) as key components of these LLM-based prompt rubrics.
We used GPT-4o-mini (OpenAI, 2024) with a Therapist-Guided Usefulness Rubric (Appendix D) to score query–document pairs sharing the same subtopic on:
@@
    Practical Helpfulness & Engagement (1–50): clarity, feasibility, motivation, and actionable guidance.
@@ 2.2.3 Hybrid Metric
We also conducted a sensitivity analysis by varying the coverage/usefulness weighting from fully coverage-based (100/0) to fully usefulness-based (0/100). As shown in Appendix 4, even under extreme weighting, corpus composition shifted only marginally, with at most 24.2% of targeted documents changing. Given this stability, we adopted the balanced 50/50 weighting as the most appropriate default.
@@ 2.3 Synthetic Document Generation and Corpus Augmentation
To evaluate our gap identification framework and its impact on retrieval quality, we augmented the mindline.sg knowledge base with synthetic resources. In our setting, synthetic documents were generated by language models to approximate the structure and tone of real-world psychoeducational resources. This approach allowed us to expand coverage in areas where authentic materials were sparse. Recent work, such as the ARES framework (Saad-Falcon et al., 2024), has even relied entirely on synthetic data for RAG evaluation. In contrast, we use synthetic documents to complement the licensed and curated materials in mindline.sg, enabling substantial expansion of the corpus and the construction of alternative corpora for testing and comparison.
@@
Using this metadata, we generated 7,424 synthetic articles with GPT-4o-mini, guided by prompts emphasizing therapist-informed principles such as emotional safety, harm reduction, and inclusivity (see Appendix E). Each article was created in about 16.1 seconds at a cost of \$0.0007.
@@ 2.3.1 Construction of Directed and Non-Directed Corpora
To assess the effect of added documents on retrieval quality, we built ten Directed corpora (gap-informed) and ten Non-Directed corpora (randomly sampled), ranging from 437 to 3,341 documents (+12.9% to +763.3% over the 387-document mindline.sg baseline). For comparison, the full Reference Corpus (8,027 documents) represents a +1,974.2% increase. All corpora include the mindline.sg baseline.
@@ 2.4.1 Retrieval Pipelines
To ensure robustness across retrieval methods, we evaluated performance using four RAG pipelines of increasing sophistication: Baseline (Lewis et al., 2020), Hierarchical (Zhang et al., 2024), Reranking (Abdallah et al., 2025), and Query Transformation (Ma et al., 2023).
In all pipelines, documents were embedded with OpenAI’s text-embedding-3-small model and stored in a FAISS vector index (cosine similarity) (Douze et al., 2025). Queries were embedded with the same model, and the retriever first produced a candidate set (up to 20 documents or chunks). Final results always consisted of the three top-ranked documents, which were passed downstream for generation. The key differences lay in how candidates were constructed and re-ranked:
@@
  • Query Transformation: Augmented retrieval by first rewriting each query for greater specificity and recall using a prompt from the RAG_Techniques repository (Diamant, 2025). The rewritten query was then used to retrieve 20 candidates from FAISS, which were re-ranked using the same LLM-based scoring as in the reranking pipeline. The top three were returned.
@@ 3 Results
Figure 2: Proportional allocation of Directed corpus across main topics.
@@
When aggregated at the main topic level, the top five most under-supported categories are Addiction (not drugs or alcohol), Depression, Anxiety, Family, and Racial, ethnic, or cultural concerns. Although Addiction’s subtopics do not appear in the top ten individually, they all fall within a narrow band just outside that range (#12–23). This consistency across subtopics drives Addiction’s overall ranking to #1, whereas other main topics include both high- and low-ranking subtopics that lower their averages.
@@
To interpret the results, instead of comparing LLM scores directly, we compare to the Reference corpus performance as a best case scenario. RAG experiments show that gap-informed augmentation (Directed) achieved $\sim$95% of Reference performance with far fewer documents. The smallest Directed corpus meeting this threshold required: 42% more docs for Query Transformation (162 docs), 74% for Hierarchical (288 docs), 74% for Reranking (288 docs), and 318% for Baseline (1230 docs) (Table 2).
Non-Directed corpora needed substantially more: 232% for Query Transformation (898 docs), 318% for Reranking (1230 docs), 403% for Hierarchical (1560 docs), and 763% for Baseline (2954 docs). In effect, random sampling required thousands more documents to reach the same quality.
Directed augmentation cuts content creation workload by 58.4% (Baseline), 81.5% (Hierarchical), 76.5% (Reranking), and 81.9% (Query Transformation) (Figure 3). These efficiency gains show that content gap analysis both accelerates retrieval quality improvements and reduces corpus expansion workload.
@@ 4 Discussion
These results have meaningful real-world implications, as these knowledge bases are essential buildings blocks for safe and reliable AI systems. In settings like mental health, where expert-reviewed content is costly and time-intensive to produce, a strategic approach offers a more sustainable path forward. Rather than generating thousands of documents indiscriminately, content creators can prioritize a focused, subtopic-level document quota informed by actual user needs. This approach is not only faster and cheaper, but also realistically helpful, providing actionable guidance for therapists and editorial teams under resource constraints. By adapting to both content gaps and operational constraints, targeted augmentation supports scalable and efficient corpus growth without sacrificing retrieval performance. However, we emphasize that our synthetic document generation process merely served as a proof of concept: the value proposition of most mental health knowledge bases remains that content is written by trusted experts (Perivolaris et al., 2025).
@@ 5 Limitations
Platform demographics: let’s talk primarily reflects youth-driven concerns, such as identity, academic stress, and relationships, whereas mindline.sg targets a broader adult population with needs that include parenting, financial precarity, or grief. This mismatch may bias gap detection and limit generalizability across age groups. While we prove the concept with a mental health forum, other sources of concerns expressed should be added to this analysis, including search terms, helpline conversations, population surveys, or other forums.
Reference corpus limitations: Our reference knowledge base from which we draw resources for augmentation is aggregated from multiple platforms, causing it to contain duplicative articles.
No supervised relevance labels: All retrieval and reranking methods in our pipeline rely on unsupervised similarity scores or LLM heuristics. Without ground-truth relevance annotations, the system may overlook domain-specific nuances like therapeutic framing or emotional appropriateness. Future work should involve trained experts to evaluate the utility of this framework, allowing for aligning with clinical guidelines and review procedures before deployment. More broadly, the lack of a standardized way to evaluate RAG in clinical settings is problematic (Amugongo et al., 2025).
Synthetic-only documents: Due to scraping restrictions on real mental health content, we used LLM-generated documents for augmentation. Although designed to resemble therapist-written materials, these documents were not medically reviewed and may lack clinical rigor.

### v2

# The Second Workshop on GenAI for Health: Potential, Trust, and Policy Compliance

@@ 1 Introduction
Access to reliable and timely mental health information is essential for early help-seeking and psychological support [26]. Kalckreuth et al. [15] found that 70.9% of psychiatric patients used the Internet for mental health purposes, most often to learn about disorders, medications, or services. With the growing reliance on self-help platforms, forums, chatbots, and apps [10], retrieving accurate and trusted psychoeducational resources has become increasingly important–both when accessed directly or when surfaced in chatbot responses through retrieval-augmented generation (RAG) [16].
@@
RAG is a natural language processing approach that bridges fluent generation and factual grounding by combining the semantic strengths of large language models (LLMs) with evidence from trusted knowledge bases [23, 18]. By conditioning outputs on retrieved documents rather than relying solely on pretrained parameters, RAG reduces hallucinations and produces more accurate, contextually appropriate, and trustworthy responses [4, 5, 30, 12].
@@
In mental health, knowledge bases that power retrieval systems usually contain articles, exercises, and therapeutic guides curated by trained professionals. These are often national platforms tailored to local contexts (e.g., seasonal affective disorder is absent in tropical countries like Singapore), but they are not designed for automated retrieval and are costly to expand–especially with evolving user needs. As a result, they often lack coverage across the full spectrum of user needs, both in content and in alignment with user style, language, and context [14, 21, 11]. For instance, a systematic review inventory by the Swedish Agency for Health Technology Assessment (SBU) identified over 2,000 evidence gaps in mental health from 2005–2020, highlighting severe content gaps where no systematic review exists or where evidence remains inconclusive [32]. We posit that one powerful way of characterizing such gaps is to overlay the knowledge base with an organic source of expression of user needs: gaps occur when topics are frequently raised by users in queries, surveys, or helpline conversations but are underrepresented or absent in the corpus. Ultimately, the effectiveness of information retrieval depends on the quality and completeness of the underlying corpus [19, 13].
@@
Recent studies have examined how gaps between user demand and document supply affect retrieval and content generation. Abián et al. [2] showed that such misalignments can be systematically identified, but their coverage analysis relies on Wikipedia pageview metrics and lacks generalizability beyond that domain. Sinha et al. [31] introduced the metric of findability, which measures how easily relevant documents can be surfaced by real queries, distinguishing it from traditional retrievability metrics. However, findability remains a passive diagnostic that does not actively resolve coverage gaps. In contrast, Kang et al. [17] proposed CCQGen, a concept-coverage–based query generation framework that iteratively produces synthetic queries to expose uncovered concepts within individual documents. While CCQGen improves internal document representation, it does not incorporate real user demand or address under-served topics at the corpus level.
More recent work has explored related approaches that associate each document with synthetic queries to enhance retrieval. For example, Raina and Gales [27] generate question-answer pairs over atomic document units to better align user queries with content, while Yang et al. [34] apply a similar idea in healthcare by linking each document to answerable synthetic queries. These methods improve retrieval within existing corpora, whereas our work focuses on corpus-level augmentation. Specifically, we integrate demand-driven gap detection with adaptive synthetic content generation to identify under-served topics and create targeted new content to close these gaps.
@@
In this paper, we take a systematic and automated approach to improving RAG methods at the source: finding and filling content gaps rather than only optimizing retrieval or ranking. To demonstrate the utility of our framework, we apply it to a large mental health knowledge base. Our contributions are threefold:
@@
  • Knowledge base gap analysis using real-world conversations. We present a systematic framework for identifying topic-level content gaps. Using the mental health resources from mindline.sg [33], Singapore’s national digital mental health platform, as a case study, we show how real users’ posts on the let’s talk anonymous mental health forum can reveal underserved needs.
@@ 2.1 Data Collection and Preprocessing
We used two primary data sources: a mental health corpus of psychoeducational resources and a real-world source of expressed user concerns. The knowledge base consisted of 387 curated documents from mindline.sg [33]. The naturalistically expressed user queries were based on 1,223 publicly available, anonymous posts in the ”Ask a Therapist” subforum of let’s talk where users can pose questions for professional counselors, thus indicating individual needs [22]. The posts were split into 978 (80%) training queries for gap identification and 245 (20%) test queries for evaluation.
@@ 2.1.1 Topic Taxonomy and Annotation
Resources and user queries were categorized into the same list of topics to enable matching. We based our topic taxonomy on the Clinician Index of Client Concerns (CLICC), a framework used by mental health professionals to categorize concerns presented by clients [9]. To improve interpretability of these 46 CLICC topics, we expanded each topic into eight more granular subtopics, refined with therapist input, yielding 368 clinically relevant categories (see Appendix A). Using GPT-4o (see Appendix B for prompt), we labeled the 978 user queries and 387 mindline.sg resources with a single representative topic and subtopic.
@@ 2.2 Content Gap Analysis
Our proposal is that knowledge base gaps can be identified by overlaying naturalistic data indicating demand for certain content and topics. We demonstrate this with posts from a mental health forum. We applied a gap analysis framework to the mindline.sg corpus, assigning each query and document a subtopic for topic-level scoring. We defined two metrics, coverage gap and usefulness gap, which were computed and combined into an equal-weight hybrid gap.
@@ 2.2.1 Coverage Gap: TF-IDF–Inspired Scoring
This metric is inspired by the Term Frequency-Inverse Document Frequency (TF-IDF) formula [29], which measures how important a word is to a particular document in a collection by weighing a term’s local frequency against its global rarity:
@@
We adapted TF–IDF from the word–document to the subtopic–query setting by redefining its components. Here, term frequency ($TF$) measures how often a subtopic appears across user queries (local demand), while inverse document frequency ($IDF$) measures how many documents in the knowledge base cover that subtopic (global supply). Subtopics with high $TF$ but low $IDF$ indicate areas where user demand exceeds available content, quantifying coverage gaps between what users seek and what information exists.
@@
where $f_{p}(t)$ is the number of queries labeled with subtopic $t$, $df(t)$ is the number of mindline.sg documents covering $t$, $D$ is the total number of documents, $c=1$ is a smoothing constant, and $\alpha=1.5$ amplifies the effect of underrepresented topics. Thus, the score highlights subtopics with high demand but low supply, identifying knowledge gaps that limit the users’ ability to find resources.
@@ 2.2.2 Usefulness Gap: LLM-as-a-Judge
We adopt the LLM-as-a-Judge paradigm, where LLMs evaluate content quality. Prior work shows that models like GPT-4 can reach about 80% agreement with human preferences, a level similar to agreement between humans, when tested on both expert-designed questions and large-scale crowdsourced comparisons [36]. This demonstrates the scalability of LLM-as-a-Judge, though limitations such as bias and reasoning gaps remain. Recent frameworks extend this approach to RAG, focusing on Context Relevance, Answer Faithfulness, and Answer Relevance [8, 28] as key components of these LLM-based prompt rubrics.
We used GPT-4o-mini [24] with a Therapist-Guided Usefulness Rubric (Appendix D) to score query–document pairs sharing the same subtopic on:
@@
    Practical Helpfulness & Engagement (1–50): clarity, feasibility, and actionable guidance.
@@ 2.2.3 Hybrid Metric
We also conducted a sensitivity analysis varying coverage/usefulness weighting from fully coverage-based (100/0) to fully usefulness-based (0/100). Corpus composition of targeted documents differed only up to 24.2% (see Appendix 5). Given this stability, we adopted a balance 50/50 weighting.
@@ 2.3 Synthetic Document Generation and Corpus Augmentation
To evaluate our gap identification framework and its impact on retrieval quality, we augmented the mindline.sg knowledge base with synthetic resources. In our setting, synthetic documents were generated by language models to approximate the structure and tone of real-world psychoeducational resources. This approach allowed us to expand coverage in areas where authentic materials were sparse. Recent work, such as the ARES framework [28], has even relied entirely on synthetic data for RAG evaluation. In contrast, we use synthetic documents to complement the licensed and curated materials in mindline.sg, enabling substantial expansion of the corpus and the construction of alternative corpora for testing and comparison.
@@
Using this metadata, we generated 7,424 synthetic articles with GPT-4o-mini, guided by prompts emphasizing therapist-informed principles such as emotional safety, harm reduction, and inclusivity (see Appendix F). Each article was created in 16.1 seconds on average at a cost of \$0.0007.
@@ 2.3.1 Construction of Directed and Non-Directed Corpora
To evaluate the impact of added documents, we built ten Directed corpora (gap-informed) and ten Non-Directed corpora (randomly sampled), ranging from 437 to 3,341 documents (+12.9% to +763.3% over the 387-document mindline.sg baseline). For comparison, the full Reference Corpus (8,027 documents) represents a +1,974.2% increase. All corpora include the mindline.sg baseline.
@@ 2.4.1 Retrieval Pipelines
To ensure robustness across retrieval methods, we evaluated performance using four RAG pipelines: Baseline [18], Hierarchical [35], Reranking [1], and Query Transformation [20].
In all pipelines, documents were embedded with OpenAI’s text-embedding-3-small model and stored in a FAISS vector index (cosine similarity) [7]. Queries were embedded with the same model, and the retriever first produced a candidate set (up to 20 documents or chunks). Final results always consisted of the three top-ranked documents, which were passed downstream for generation. The key differences lay in how candidates were constructed and re-ranked:
@@
  • Query Transformation: Augmented retrieval by first rewriting each query for greater specificity and recall using a prompt from the RAG_Techniques repository [6]. The rewritten query was then used to retrieve 20 candidates from FAISS, which were re-ranked using the same LLM-based scoring as in the reranking pipeline. The top three were returned.
@@ 3 Results
Figure 2: Proportional distribution of Directed corpus across main topics.
@@
When aggregated at the main topic level, the top five most under-supported categories are Addiction (not drugs or alcohol), Depression, Anxiety, Family, and Racial, ethnic, or cultural concerns. Although Addiction’s subtopics do not rank in the top ten individually, they cluster just below that range (#12–23), elevating Addiction to #1 overall while other topics combine high- and low-ranking subtopics that lower their averages.
@@
To interpret the results, instead of comparing LLM scores directly, we compare to the Reference corpus performance as a best case scenario. RAG experiments show that gap-informed augmentation (Directed) achieved $\sim$95% of Reference performance with far fewer documents. The smallest Directed corpus meeting this threshold required: 42% more docs for Query Transformation (162 docs), 74% for Hierarchical (288 docs), 74% for Reranking (288 docs), and 318% for Baseline (1,230 docs) (Table 2). Non-Directed corpora needed more: 232% for Query Transformation (898 docs), 318% for Reranking (1,230 docs), 403% for Hierarchical (1,560 docs), and 763% for Baseline (2,954 docs). Random sampling therefore required thousands more documents to reach the same quality. Thus, Directed augmentation cuts content creation workload by 58.4% (Baseline), 81.5% (Hierarchical), 76.5% (Reranking), and 81.9% (Query Transformation) (Figure 3), demonstrating that gap analysis accelerates retrieval gains while minimizing corpus expansion effort.
@@ 4 Discussion
These results have meaningful real-world implications, as these knowledge bases are essential building blocks for safe and reliable AI systems. In settings like mental health, where expert-reviewed content is costly and time-intensive to produce, a strategic approach offers a more sustainable path forward. Rather than generating thousands of documents indiscriminately, content creators can prioritize a focused, subtopic-level document quota informed by actual user needs. This approach is not only faster and cheaper, but also realistically helpful, providing actionable guidance for therapists and editorial teams under resource constraints. By adapting to both content gaps and operational constraints, targeted augmentation supports scalable and efficient corpus growth without sacrificing retrieval performance. However, we emphasize that our synthetic document generation process merely served as a proof of concept: the value proposition of most mental health knowledge bases remains that content is written by trusted experts [25].
While gap-informed augmentation improved overall retrieval, some niche or emergent queries still performed poorly due to limited source material. Examples include relationship insecurity influenced by social media use, ongoing career or academic setbacks associated with low self-esteem, and expressions of emotional detachment or confusion about empathy. These cases indicate where human content creators should focus on new article development.
@@ 5 Limitations
Platform demographics mismatch: let’s talk primarily reflects youth-driven concerns, such as identity, academic stress, and relationships, whereas mindline.sg targets a broader adult population with needs that include parenting, financial precarity, or grief. Including additional sources of expressed concerns can mitigate bias in gap detection and improve generalizability across demographic groups.
Reference corpus limitations: Our reference knowledge base aggregates content from multiple mental health platforms, leading to some duplication. While pruning could reduce redundancy, we note that partial overlap is typical in mental health platforms, where similar psychoeducational topics are intentionally reiterated across sites to maintain accessibility and consistency of guidance. Future work could examine pruning or deduplication to quantify its marginal impact on retrieval quality.
No supervised relevance labels: All retrieval and reranking methods rely on unsupervised similarity scores or LLM-as-a-Judge evaluations without human or clinical validation. Without ground-truth relevance annotations, the system may overlook domain-specific nuances like therapeutic framing or emotional appropriateness. Future work should involve trained experts to evaluate the utility of this framework, allowing for aligning with clinical guidelines and review procedures before deployment. More broadly, the lack of a standardized way to evaluate RAG in clinical settings is problematic [3].
Synthetic-only documents: Due to scraping restrictions on real mental health content, we used LLM-generated documents for augmentation. While modeled after therapist-written materials, they were not medically reviewed and served only as a proof of concept. In practice, all new content should be authored by qualified professionals to ensure clinical accuracy and maintain user trust.